OpenStudy (anonymous):

How to simplify : (1 - 1/(2^2)).(1 - 1/(3^2)).(1 - 1/(4^2)).(1 - 1/(5^2))...(1 - 1/(2010^2))

OpenStudy (anonymous):

is it ((2-1)(2+1))(3-1)(3+1)(4-1)(4+1)(5-1)(5+1)........(2010-1)(2010+1)/(2.3.4.5...2010)^2 = (2-1)*(2010+1)/(2*2010)=0.50025 appx

OpenStudy (anonymous):

OK, thanks...